Selected information regarding a companys most recent quarter follows (all data in thousands) Ending work in process inventory Cost of goods manufactured Direct labor Direct materials used Beginning work in process inventory S650 S830 S430 $170 $400 What was manufacturing overhead for the quarter? OA $580 OB. S230 Ос. s480 O D. S600
Answer #1

Cost of goods manufactured=Total manufacturing cost+Beginning WIP-Ending WIP

Total manufacturing cost=(830+650-400)=$1080

Total manufacturing cost=Direct material used+Direct labor+Manufacturing overhead

Manufacturing overhead=(1080-170-430)

=$480